American Health Systems has 6,500,000 shares of stock outstanding and will report earnings of $12 million in the current year. The company is considering the issuance of 1,300,000 additional shares, which can only be issued at $30 per share.


a. Assume that American Health Systems can earn 7 percent on the proceeds. Calculate earnings per share.

b. Should the new issue be undertaken based on earnings per share? Yes or No
Answer #1

a.

Amount received on new issue =1300000*30 = 39,000,000

earnings on 39,000,000 @ 7% = 2,730,000

New EPS =(12,000,000+2,730,000)/(6,500,000+1,300,000) =1.89

b.

Current EPS was =12000000/6500000 =1.85

EPS has increased, so the issue should be undertaken.